P1= t1p2/t2 solve for t2
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

P1= t1p2/t2
P1*t2 = t1*p2
t2 = t1*p2/P1

Answer by solver91311(24713)





Multiply both sides by and then simplify.

John

My calculator said it, I believe it, that settles it